KINETICS MEASUREMENTS OF HO_{2} AND RO_{2} SELF AND CROSS REACTIONS USING INFRARED KINETIC SPECTROSCOPY (IRKS)

\underline{A. C. NOELL}, L. S. ALCONCEL, D. J. ROBICHAUD, and M. OKUMURA, Department of Chemistry, California Institute of Technology, Pasadena, CA 91125; S. P. SANDER, Jet Propulsion Laboratory, Pasadena, CA 91109.

Reactions between RO_{2} and HO_{2} are important in the gas phase oxidation of volatile organics in the atmosphere. The rate constants for the reactions of HO_{2} + C_{2}H_{5}O_{2}, and C_{2}H_{5}O_{2} + C_{2}H_{5}O_{2} were measured using a novel approach to radical radical reactions that employs two independent spectroscopic probes. HO_{2} was monitored at 6638.2 cm^{−1} using near-IR diode laser WM spectroscopy, and C_{2}H_{5}O_{2} was monitored at 250 nm using UV absorption. Experiments were run under conditions relevant to the upper troposphere (221-298 K and 50-400 Torr). Measurements of the rate constants for both reactions as well as a measurement of the branching fraction leading to the alkoxy channel of the C_{2}H_{5}O_{2} self reaction will be discussed. Preliminary results are: k_{HC} = (6.4 ± 1.4) × 10^{−13} cm^{3} molec^{−1} s^{−1} Exp((601 ± 120)/ T ), k_{CC} = (1.10 ± 0.20) × 10^{−13} cm^{3} molec^{−1} s^{−1} independent of temperature, and a branching fraction of 0.32 ± 0.13.